OpenStudy (anonymous):

Solve using variation of parameters. \[y''-4y=\frac{e^{2x}}{x}\] \[y_1=e^{2x}, y_2=e^{-2x}, y_p=u_1y_1+u_2y_2\] \[W = \begin{vmatrix} e^{2x} & e^{-2x}\\2e^{2x} & -2e^{-2x} \end{vmatrix} = -4\] \[W_1 = \begin{vmatrix} 0 & e^{-2x}\\ \frac{e^{2x}}{x} & -2e^{-2x} \end{vmatrix} = -x^{-1}\] \[W_2 = \begin{vmatrix} e^{2x} & 0\\2e^{2x} & \frac{e^{2x}}{x} \end{vmatrix} = \frac{e^{4x}}{x}\] \[u'_1 = \frac{W_1}{W}=\frac{1}{4x}, u_1=\frac{1}{4}\ln x\] \[u'_2 = \frac{W_2}{W}=-\frac{e^{4x}}{4x},u_2=-\int \frac{e^{4x}}{4x} dx =?\]

OpenStudy (phebe):



Miracrown (miracrown):

Alrighty so this is differential equations. I haven't taken this in years, however, I was good at it when I took it. I think I might be able to help. I would just need to look up variation of parameters in my book.

Miracrown (miracrown):

oh the Wronskian

zepdrix (zepdrix):

Hmm I can't find any mistakes in your work.... That last integral is the `exponential integral`: noted by \(\Large\rm Ei(x)\) It's not solvable using normal methods :o Maybe uhhh power series or something... hmm thinking

Miracrown (miracrown):

so we need to integrate that

Miracrown (miracrown):

hm, the difficult part is the 1/4x part

Miracrown (miracrown):

because we know how to integrate e^4x

Miracrown (miracrown):

Have you tried integration by parts ? @heril @zepdrix yep, looks like you're right hmm...

zepdrix (zepdrix):

\[\Large\rm -\frac{1}{4}\int\limits \frac{e^{4x}}{x} dx =-\frac{1}{4}\int\limits \frac{1+4x+\frac{(4x)^2}{2!}+\frac{(4x)^3}{3!}+...}{x} dx\]You could write the power series for e^4x like that, and then divide an x out of each term. The first term ends up giving you ln(x) and then I bet after we integrate we can group the other terms somehow maybe :o hmmm

Miracrown (miracrown):

 yep, not seeing an easy way to do that except specifying that it is the Expontial integral